# Multi-stage tomography based on eigenanalysis for high-dimensional dense unitary quantum processes

### v1

@@ II.2 Second part of the methods
What is available in practice is a ket $\lvert\widehat{\Psi}_{2}\rangle$, obtained by performing a QST at the output of the process, and equal to $\lvert\Psi_{2}\rangle$ up to an unknown phase factor and to estimation errors. If the considered QST algorithm does not guarantee that the norm of $\lvert\widehat{\Psi}_{2}\rangle$ is equal to 1, a recommended pre-processing step consists of dividing $\lvert\widehat{\Psi}_{2}\rangle$ by its norm. When neglecting the above estimation errors,
@@
By the way, based on the above condition on nonzero components for $\lvert\Psi_{1}\rangle$, we preferably select a state $\lvert\Psi_{1}\rangle$ with all component moduli as high as possible. Considering positive real-valued components and taking into account that $\lvert\Psi_{1}\rangle$ has unit norm, this yields a single solution: all components of $\lvert\Psi_{1}\rangle$ in the considered basis are equal to $1/\sqrt{d}$.
Going back to (14), the associated matrix form reads
@@
i.e. $\widehat{U}_{5}$ succeeds in restoring $U$, up to a global phase factor, that has no physical consequences as explained above (and again up to the estimation errors that were mentioned above but implicit in the above equations).
@@
Input : a) Estimate $\widehat{\rho}_{2}$ of output density matrix $\rho_{2}$ (provided by Quantum State Tomography (QST) and obtained for input density matrix (8)). b) Estimate $\lvert\widehat{\Psi}_{2}\rangle$ of output ket $\lvert\Psi_{2}\rangle$ (provided by QST and obtained for all components of input ket $\lvert\Psi_{1}\rangle$ equal to $1/\sqrt{d}$).
@@
       /* [see Sect. II.1] If this is not performed by the considered QST algorithm, then derive Hermitian and unit-trace version $\widehat{\rho}_{4}$ of $\widehat{\rho}_{2}$ as follows (otherwise, $\widehat{\rho}_{4}=\widehat{\rho}_{2}$): */
       Compute $\widehat{\rho}_{3}$ and $\widehat{\rho}_{4}$ by using (2) and (3);
@@
       /* [see Sect. II.2] If the QST algorithm does not yield a unit-norm ket $\lvert\widehat{\Psi}_{2}\rangle$, then create it as follows: */
       Divide $\lvert\widehat{\Psi}_{2}\rangle$ by its norm;
@@
       Compute $\lvert\Psi_{3}\rangle$ by using (12);
       Compute $\widehat{U}_{5}$ by using (16);
@@
Algorithm 1 EQPT1: first Eigenanalysis-based Quantum Process Tomography algorithm (composed of one stage).
@@ III.1 Motivation
III.2 First part of the methods
As in Section II.1, the method proposed here starts by considering a given mixed state that has a diagonal density matrix $\rho_{1}$ and that is applied to the input of the considered unitary process defined by a matrix $U$. However, we here accept to have identical values on the main diagonal of $\rho_{1}$, at the expense of then having to handle more complex eigendecomposition properties for the process output, as shown below. More precisely, each value that appears on the main diagonal of $\rho_{1}$ appears $d_{1}$ times, on adjacent rows, and that diagonal contains $d_{2}$ different values. The number of rows of $\rho_{1}$, equal to the dimension of the considered state space, is thus equal to
@@ III.2 First part of the methods
(in practice, we aim at considering the case when $d_{1}$, $d_{2}$, and hence $d$ are powers of 2, with an application to qubits). If requesting the same minimum gap between different eigenvalues as in Section II.1, i.e. replacing the bound (18) on $d$ in Section II by the same bound value for $d_{2}$ instead of $d$ here, the maximum value of the state space dimension (19) that can be handled here is $d_{1}$ times higher than the maximum dimension that can be addressed by the methods of Section II, which is our goal, ideally. However, it is not guaranteed that the methods proposed hereafter can handle the same gap between different eigenvalues as in Section II.1 with the same accuracy, especially due to the more complex structure of these extended methods. The practical accuracy of all proposed methods is therefore investigated numerically further in this paper.
@@
with each diagonal value repeated $d_{1}$ times. Without loss of generality, we here again request the values on the diagonal of $\rho_{1}$ to be in nonincreasing order i.e. $r_{1}>r_{2}>\dots>r_{d_{2}}$. Moreover, here again preferably using a uniform spacing between these values, (8) is here replaced by:
@@
Here again, the process output corresponding to $\rho_{1}$ is a mixed state and its density matrix $\rho_{2}$ may be expressed as (1). Moreover, as in Section II.1, $\rho_{2}$ is Hermitian and hence unitarily diagonalizable, and it has unit trace. Besides, in practice, an estimate $\widehat{\rho}_{2}$ of $\rho_{2}$ is derived by a QST algorithm and, if required by the considered QST algorithm, we preprocess this estimate of $\rho_{2}$ in order to derive its Hermitian and unit-trace version, denoted as $\widehat{\rho}_{4}$. We then perform an eigendecomposition of $\widehat{\rho}_{4}$. The analysis of the properties of this eigendecomposition here requires additional care, because the considered matrix has various identical eigenvalues.
@@
The properties of Section II.1 that remain true here are as follows. The eigendecomposition is not unique, Eq. (4)-(5) still apply, $V_{4}$ and $D_{4}$ are respectively a unitary matrix and a diagonal matrix, the $k$th column of $V_{4}$ is an eigenvector of $\widehat{\rho}_{4}$ and the associated eigenvalue is the $k$th value on the diagonal of $D_{4}$ (see again Theorem 2 for the latter property).
In contrast, what is different here as compared with Section II.1 is that the same eigenvalue here appears $d_{1}$ time and is thus associated with $d_{1}$ eigenvectors, instead of only one in Section II.1. If first considering a completely arbitrary diagonalizable matrix, one should distinguish between the following two concepts related to eigenvalues. The algebraic multiplicity of an eigenvalue is the number of times it appears in the diagonal form (here $D_{4}$) of the considered matrix, and hence the number of corresponding first-order factors in the characteristic polynomial of that matrix (see [43] p. 58). The geometric multiplicity of an eigenvalue is the dimension of the eigenspace spanned by all the eigenvectors associated with that eigenvalue (see [43] p. 57 and p. 58). The geometric multiplicity of an eigenvalue is always lower than or equal to its algebraic multiplicity (see again [43] p. 58). We can illustrate as follows the case that yields the highest difference between the geometric and algebraic multiplicities. Still considering arbitrary matrices, i.e. not yet focusing on the more specific scope of the present paper, we can choose a unitary matrix $V_{4}$ and a diagonal matrix $D_{4}$, and we can create the matrix $\widehat{\rho}_{4}$ defined by (4). Moreover, we can choose $D_{4}$ with the desired high algebraic multiplicity for one (or several) of its eigenvalues and we can use the same value for all eigenvectors of $V_{4}$ associated with that repeated eigenvalue. Then, the geometric multiplicity of that eigenvalue is equal to its lowest possible value, i.e. equal to one, hence (possibly much) lower than its algebraic multiplicity. We again stress that the above properties are for a completely arbitrary (diagonalizable) matrix and this shows that eigendecomposition properties should here be handled differently from what was done in Section II.1.
@@
Fortunately, we can here moreover take advantage of the fact that the matrix $\widehat{\rho}_{4}$ to be decomposed is Hermitian. For this class of matrices (and for all normal matrices), the geometric multiplicity of each eigenvalue is guaranteed to be equal to its algebraic multiplicity, i.e. these matrices are nondefective (see [43] p. 103). In other words, each eigenvalue of $\widehat{\rho}_{4}$ is here known to appear $d_{1}$ times at arbitrary locations on the diagonal of $D_{4}$ (algebraic multiplicity), since these eigenvalues are the same, in a different order, as in another eigendecomposition of the same matrix $\widehat{\rho}_{4}$ (neglecting estimation errors), namely the eigendecomposition (1). Therefore, the dimension of the eigenspace associated with any eigenvalue of $\widehat{\rho}_{4}$ is also equal to $d_{1}$. So, when performing the eigendecomposition of $\widehat{\rho}_{4}$, the considered decomposition algorithm should provide $d_{1}$ linearly independent eigenvectors for each eigenvalue, and we hereafter assume that practical eigendecomposition algorithms succeed in finding them (that was indeed the case in all the tests reported further in this paper). As an overall result, when performing the eigendecomposition of $\widehat{\rho}_{4}$, 1) we get a diagonal matrix $D_{4}$ with $d=d_{1}d_{2}$ eigenvalues that may be in an arbitrary order, and with each encountered value repeated $d_{1}$ times and 2) we get a unitary matrix $V_{4}$ whose columns are eigenvectors that have the above-defined properties.
@@
As in Section II.1, we then apply a joint permutation to the diagonal elements of $D_{4}$ and to the columns of $V_{4}$, so that the values corresponding to the diagonal elements of $D_{4}$ become in nonincreasing order. Considering the resulting complete set of values from the top to the bottom of the diagonal of the reordered version of $D_{4}$, we thus get $d_{2}$ successive subsets, with the same value repeated $d_{1}$ times in each subset (in practice, these $d_{1}$ values may be slightly different, due to estimation errors), and with different and decreasing values from one subset to the next one. In other words, up to estimation errors, the reordered version of $D_{4}$ is nothing but the matrix $\rho_{1}$ of (20). Each subset of identical values along the diagonal of the obtained matrix (values equal to eigenvalue $r_{m_{1}}$) has an index $m_{1}$, with $m_{1}\in\{1,\dots,d_{2}\}$, and consists of elements with indices  [^3] In this whole paper except Appendix F, the index of the first row and column of a matrix is equal to 1.
@@
III.3 Second part of the methods
The two important properties that we showed in Section III.2 for the considered (possibly high) dimension $d=d_{1}d_{2}$ are as follows. First, we proved that we can “partition” the set of $d$ columns of $U$ in $d_{2}$ subsets, in the sense that, in Section III.2, we identified each $d_{1}$-dimensional subspace ${\cal S}_{1,m_{1}}$ spanned by the columns of $U$ that compose the subset with index $m_{1}$. Second, we proved that we can select which partition we perform: the columns of $U$ that are gathered in the same subset are those that correspond to the same eigenvalue of $\rho_{2}$ and hence $\widehat{\rho}_{4}$.
In the present section, we take advantage of these results to create another partition of the columns of $U$, in order to then jointly exploit these two partitions as explained further in Section III.4. We thus create a two-stage algorithm in the sense that it successively uses two eigendecompositions of density matrices estimated at the output of the process $U$, namely the eigendecompositions respectively described in Section III.2 and in the present subsection. This should be contrasted with the algorithms of Section II, that are single-stage methods in the sense that they perform a single eigendecomposition.
More precisely, the input density matrix $\rho_{5}$ here applied to the considered unitary process is different from $\rho_{1}$ but, here again, it is diagonal, it contains $d_{2}$ different values that each appear $d_{1}$ times and these values are the same as in Section III.2. However, they are here placed in a different order: the set of values on the main diagonal of $\rho_{5}$ consists of a series of $d_{1}$ identical subsets, each of which successively contains the $d_{2}$ values $r_{1},\dots,r_{d_{2}}$. The density matrix $\rho_{5}$ thus reads
@@ III.3 Second part of the methods
This input density matrix $\rho_{5}$ is used in the same way as in Section III.2. So, the density matrix $\rho_{6}$ of the corresponding process output may be derived from $\rho_{5}$ in the same way as in (1). An estimate $\widehat{\rho}_{6}$ of $\rho_{6}$ is derived by a QST algorithm and, if required by the considered QST algorithm, we preprocess $\widehat{\rho}_{6}$ in order to derive: 1) its Hermitian version $\widehat{\rho}_{7}$ (in the same way as in (2)) and then 2) its Hermitian and unit-trace version $\widehat{\rho}_{8}$ (in the same way as in (3)). We then perform an eigendecomposition of $\widehat{\rho}_{8}$. This yields a diagonal matrix $D_{8}$ containing its eigenvalues and a matrix $V_{8}$ whose columns are its associated eigenvectors.
@@
Moreover, the corresponding matrix, derived from $V_{8}$, with reordered eigenvectors in its columns is denoted as $\widehat{U}_{3}$  [^4] In this whole paper, all the quantities of importance related to estimates of the considered process are denoted with the same letter and increasing indices 1, 2 and so on. When describing our first type of methods in Section II, for the sake of clarity we used detailed notations for these estimates, i.e. with a first notation (with index 1) for the quantity before column reordering and a second notation (with index 2) for the quantity after column reordering. However, we then did not use the first notation. We here want to make it clear that, for the second type of methods described in the present section, we also consider two such quantities, i.e. before and after reordering. However, we do not introduce a notation for the quantity before reordering because, here again, we would not use it afterwards. Therefore, we here only introduce a single quantity, with an index equal to 3.. Its subset of columns that also have the indices defined by (24) is a set of $d_{1}$ eigenvectors that form a basis of the $d_{1}$-dimensional eigenspace of $\widehat{\rho}_{8}$ associated with its eigenvalue $r_{m_{2}}$. This eigenspace, obtained in the second part of this method, is hereafter denoted as ${\cal S}_{2,m_{2}}$.
${\cal S}_{2,m_{2}}$ also has a relationship with the actual matrix $U$ to be identified, but it is different from the relationship disclosed for ${\cal S}_{1,m_{1}}$ in Section III.2. Here again, ${\cal S}_{2,m_{2}}$ is the $d_{1}$-dimensional subspace spanned by the $d_{1}$ columns of $U$ that correspond to the eigenvalue $r_{m_{2}}$. However, here the indices of these columns are different from those of Section III.2: they are equal to the column indices of all occurrences of $r_{m_{2}}$ in (23), that is
@@
III.4 Third part of the methods

### v2

@@ II-B Second part of the methods
What is available in practice is a ket $\lvert\widehat{\Psi}_{2}\rangle$, obtained by performing a QST at the output of the process, and equal to $\lvert\Psi_{2}\rangle$ up to an unknown phase factor and to estimation errors (see again Appendix C concerning a possible normalization of the QST output used here). When neglecting the above estimation errors,
@@
The associated matrix form reads
@@
which means that $\widehat{U}_{5}$ succeeds in restoring $U$, up to a global phase factor (and again up to the estimation errors that were mentioned above but implicit in the above equations).
@@
Input : a) Estimate $\widehat{\rho}_{4}$ of output density matrix $\rho_{2}$ (provided by Quantum State Tomography (QST) and obtained for input density matrix (40)). b) Estimate $\lvert\widehat{\Psi}_{2}\rangle$ of output ket $\lvert\Psi_{2}\rangle$ (provided by QST and obtained for all components of input ket $\lvert\Psi_{1}\rangle$ equal to $1/\sqrt{d}$).
@@
       $\lvert\Psi_{3}\rangle=\widehat{U}_{2}^{{\dagger}}\lvert\widehat{\Psi}_{2}\rangle$;
       $\widehat{U}_{5}=\widehat{U}_{2}\hskip 2.84526pt\mathrm{diag}(\lvert\Psi_{3}\rangle\oslash\lvert\Psi_{1}\rangle)$;
@@
Algorithm 1 EQPT1: first Eigenanalysis-based Quantum Process Tomography algorithm (composed of one stage). This version is for the case when the QST algorithm provides estimates that meet the known properties of the actual quantities. Otherwise, see Appendix C.
@@ III-A Motivation
III-B First part of the methods
As in Section II-A, the method proposed here starts by considering a given mixed state that has a diagonal density matrix $\rho_{1}$ and that is applied to the input of the considered unitary process defined by a matrix $U$. However, we here accept to have identical values on the main diagonal of $\rho_{1}$, at the expense of then having to handle more complex eigendecomposition properties for the process output, as shown below. More precisely, each value that appears on the main diagonal of $\rho_{1}$ appears $d_{1}$ times, on adjacent rows, and that diagonal contains $d_{2}$ different values. The number of rows of $\rho_{1}$, equal to the dimension of the considered state space, is thus equal to
@@ III-B First part of the methods
(in practice, we aim at considering the case when $d_{1}$, $d_{2}$, and hence $d$ are powers of 2, with an application to qubits). If requesting the same minimum gap between different eigenvalues as in Section II-A, or in other words, if replacing the bound (15) on $d$ in Section II by the same bound value for $d_{2}$ instead of $d$ here, then the maximum value of the state space dimension (16) that can be handled here is $d_{1}$ times higher than the maximum dimension that can be addressed by the methods of Section II, which is our goal, ideally. However, it is not guaranteed that the methods proposed hereafter can handle the same gap between different eigenvalues as in Section II-A with the same accuracy, especially due to the more complex structure of these extended methods. The practical accuracy of all proposed methods is therefore investigated numerically further in this paper.
@@
with each diagonal value repeated $d_{1}$ times. This may also be expressed as
$\rho_{1}=\mathrm{diag}(r)\otimes I_{d_{1}}$ (18)
where $r$ is the vector that contains the values $r_{1},\dots,r_{d_{2}}$, whereas $I_{d_{1}}$ is the $d_{1}\times d_{1}$ identity matrix and $\otimes$ stands for the Kronecker product [53]. Without loss of generality, we here again request the values on the diagonal of $\rho_{1}$ to be in nonincreasing order: $r_{1}>r_{2}>\dots>r_{d_{2}}$. Moreover, we preferably use values that are uniformly distributed (that is, multiples of the same step, with zero excluded), as in the method of Section II (see details for the latter method in Appendix D). In the configuration of the present section, (40) is thus replaced by:
@@
Here again, the process output corresponding to $\rho_{1}$ is a mixed state and its density matrix $\rho_{2}$ may be expressed as (1). Moreover, as in Section II-A, $\rho_{2}$ is Hermitian and hence unitarily diagonalizable, and it has unit trace. Besides, in practice, an estimate $\widehat{\rho}_{2}$ of $\rho_{2}$ is derived by a QST algorithm and, if required by the considered QST algorithm, we preprocess this estimate of $\rho_{2}$ in order to derive its Hermitian and unit-trace version, denoted as $\widehat{\rho}_{4}$. We then perform an eigendecomposition of $\widehat{\rho}_{4}$. The analysis of the properties of this eigendecomposition here requires additional care, because the considered matrix has various identical eigenvalues.
@@
The properties of Section II-A that remain true here are as follows. The eigendecomposition is not unique, Eq. (2)-(3) still apply, $V_{4}$ and $D_{4}$ are respectively a unitary matrix and a diagonal matrix, the $k$th column of $V_{4}$ is an eigenvector of $\widehat{\rho}_{4}$ and the associated eigenvalue is the $k$th value on the diagonal of $D_{4}$ (see again Theorem 2 for the latter property).
@@
In contrast, what is different here as compared with Section II-A is that the same eigenvalue here appears $d_{1}$ time and is thus associated with $d_{1}$ eigenvectors, instead of only one in Section II-A. Eigendecomposition properties are more complex than above in this general framework but, fortunately, their complexity is here limited by the fact that the matrix $\widehat{\rho}_{4}$ to be decomposed is Hermitian. For this class of matrices (and for all normal matrices), the so-called geometric multiplicity of each eigenvalue is guaranteed to be equal to its so-called algebraic multiplicity, i.e. these matrices are nondefective (see details in [48] p. 57, 58, 103). More explicitly, this means that each eigenvalue of $\widehat{\rho}_{4}$ is here known to appear $d_{1}$ times at arbitrary locations on the diagonal of $D_{4}$ (algebraic multiplicity), since these eigenvalues are the same, in a different order, as in another eigendecomposition of the same matrix $\widehat{\rho}_{4}$ (neglecting estimation errors), namely the eigendecomposition (1). Therefore, the dimension of the eigenspace associated with any eigenvalue of $\widehat{\rho}_{4}$ is also equal to $d_{1}$. So, when performing the eigendecomposition of $\widehat{\rho}_{4}$, the considered decomposition algorithm should provide $d_{1}$ linearly independent eigenvectors for each eigenvalue, and we hereafter assume that practical eigendecomposition algorithms succeed in finding them (that was indeed the case in all the tests reported further in this paper). As an overall result, when performing the eigendecomposition of $\widehat{\rho}_{4}$, 1) we get a diagonal matrix $D_{4}$ with $d=d_{1}d_{2}$ eigenvalues that may be in an arbitrary order, and with each encountered value repeated $d_{1}$ times and 2) we get a unitary matrix $V_{4}$ whose columns are eigenvectors that have the above-defined properties.
@@
As in Section II-A, we then apply a joint permutation to the diagonal elements of $D_{4}$ and to the columns of $V_{4}$, so that the values corresponding to the diagonal elements of $D_{4}$ become in nonincreasing order. Considering the resulting complete set of values from the top to the bottom of the diagonal of the reordered version of $D_{4}$, we thus get $d_{2}$ successive subsets, with the same value repeated $d_{1}$ times in each subset (in practice, these $d_{1}$ values may be slightly different, due to estimation errors), and with different and decreasing values from one subset to the next one. In other words, up to estimation errors, the reordered version of $D_{4}$ is nothing but the matrix $\rho_{1}$ of (17). Each subset of identical values along the diagonal of the obtained matrix (values equal to eigenvalue $r_{m_{1}}$) has an index $m_{1}$, with $m_{1}\in\{1,\dots,d_{2}\}$, and consists of elements with indices [^3] In this whole paper except Appendix K, the index of the first row and column of a matrix is equal to 1.
@@
III-C Second part of the methods
The two important properties that we showed in Section III-B for the considered (possibly high) dimension $d=d_{1}d_{2}$ are as follows. First, we proved that we can “partition” the set of $d$ columns of $U$ in $d_{2}$ subsets, in the sense that, in Section III-B, we identified each $d_{1}$-dimensional subspace ${\cal S}_{1,m_{1}}$ spanned by the columns of $U$ that compose the subset with index $m_{1}$. Second, we proved that we can select which partition we perform: the columns of $U$ that are gathered in the same subset are those that correspond to the same eigenvalue of $\rho_{2}$ and hence $\widehat{\rho}_{4}$.
In the present section, we take advantage of these results to create another partition of the columns of $U$, in order to then jointly exploit these two partitions as explained further in Section III-D. We thus create a two-stage algorithm in the sense that it successively uses two eigendecompositions of density matrices estimated at the output of the process $U$, namely the eigendecompositions respectively described in Section III-B and in the present subsection. This should be contrasted with the algorithms of Section II, that are single-stage methods in the sense that they perform a single eigendecomposition.
More precisely, the input density matrix $\rho_{5}$ here applied to the considered unitary process is different from $\rho_{1}$ but, here again, it is diagonal, it contains $d_{2}$ different values that each appear $d_{1}$ times and these values are the same as in Section III-B. However, they are here placed in a different order: the set of values on the main diagonal of $\rho_{5}$ consists of a series of $d_{1}$ identical subsets, each of which successively contains the $d_{2}$ values $r_{1},\dots,r_{d_{2}}$. The density matrix $\rho_{5}$ thus reads
@@ III-C Second part of the methods
This may also be expressed as
$\rho_{5}=I_{d_{1}}\otimes\mathrm{diag}(r)$ (22)
with the same notations as in (18).
This input density matrix $\rho_{5}$ is used in the same way as in Section III-B. So, the density matrix $\rho_{6}$ of the corresponding process output may be derived from $\rho_{5}$ in the same way as in (1). An estimate $\widehat{\rho}_{6}$ of $\rho_{6}$ is derived by a QST algorithm and, if required by the considered QST algorithm, we preprocess $\widehat{\rho}_{6}$ in order to derive: 1) its Hermitian version $\widehat{\rho}_{7}$ (in the same way as in (38)) and then 2) its Hermitian and unit-trace version $\widehat{\rho}_{8}$ (in the same way as in (39)). We then perform an eigendecomposition of $\widehat{\rho}_{8}$. This yields a diagonal matrix $D_{8}$ containing its eigenvalues and a matrix $V_{8}$ whose columns are its associated eigenvectors.
@@
Moreover, the corresponding matrix, derived from $V_{8}$, with reordered eigenvectors in its columns is denoted as $\widehat{U}_{3}$  [^4] In this whole paper, all the quantities of importance related to estimates of the considered process are denoted with the same letter and increasing indices 1, 2 and so on. When describing our first type of methods in Section II, for the sake of clarity we used detailed notations for these estimates, therefore with a first notation (with index 1) for the quantity before column reordering and a second notation (with index 2) for the quantity after column reordering. However, we then did not use the first notation. We here want to make it clear that, for the second type of methods described in the present section, we also consider two such quantities, namely before and after reordering. However, we do not introduce a notation for the quantity before reordering because, here again, we would not use it afterwards. Therefore, we here only introduce a single quantity, with an index equal to 3.. Its subset of columns that also have the indices defined by (23) is a set of $d_{1}$ eigenvectors that form a basis of the $d_{1}$-dimensional eigenspace of $\widehat{\rho}_{8}$ associated with its eigenvalue $r_{m_{2}}$. This eigenspace, obtained in the second part of this method, is hereafter denoted as ${\cal S}_{2,m_{2}}$.
${\cal S}_{2,m_{2}}$ also has a relationship with the actual matrix $U$ to be identified, but it is different from the relationship disclosed for ${\cal S}_{1,m_{1}}$ in Section III-B. Here again, ${\cal S}_{2,m_{2}}$ is the $d_{1}$-dimensional subspace spanned by the $d_{1}$ columns of $U$ that correspond to the eigenvalue $r_{m_{2}}$. However, here the indices of these columns are different from those of Section III-B: they are equal to the column indices of all occurrences of $r_{m_{2}}$ in (21), that is
@@
III-D Third part of the methods